# Unified description of sum rules and duality between CP phases and unitarity triangles through third-order rephasing invariants

### v1

@@ Abstract
In this letter, we demonstrate that products of third-order rephasing invariants $V_{\alpha i}V_{\beta j}V_{\gamma k}/\det V$ of flavor mixing matrix $V$ reproduce all the nine angles of unitarity triangles and all the CP phases in the nine parameterizations of $V$. The sum rules relating the CP phases and angles are also decomposed into terms of these rephasing invariants. Furthermore, through ninth-order invariants, these fourth- and fifth-order invariants become equivalent, which can be regarded as a certain duality. For the phase matrix $\Delta$ and the angle matrix $\Phi$, $\Delta\pm\Phi$ are expressed in terms of even-permutations $X$ and odd-permutations $\Psi$ of third-order invariant. As a result, these are represented by the two concise matrix equations $\Phi=\Psi-{\rm X}$ and $\Delta=\Pi^{\prime}-\Psi-{\rm X}$.
@@ I Introduction
The flavor mixing matrix provides an essential formalism of flavor physics. As its CP phase is intimately related to the origin of the matter-antimatter asymmetry, it is one of the key observables for particle physics in the near future. In order to establish a theoretical basis, rephasing invariants of the mixing matrix Wu:1985ea ; Bjorken:1987tr ; Nieves:1987pp ; Jenkins:2007ip have been studied to define CP-violating quantities and to elucidate their physical significance. The fourth-order rephasing invariants, represented by the Jarlskog invariant Jarlskog:1985ht , have been widely employed in the analysis of unitarity triangles Wu:1994di ; Xing:2009eg ; Harrison:2009bz ; Antusch:2009hq ; Frampton:2010ii ; Dueck:2010fa ; Frampton:2010uq ; Li:2010ae ; Qin:2010hn ; Zhou:2011xm ; Qin:2011ub ; Zhang:2012bk ; Zhang:2012ys ; Li:2012zxa ; He:2013rba ; He:2016dco ; Xing:2019tsn ; Kaur:2023ypg ; Harrison:2025rkp . However, they are insufficient for providing a unified understanding of the CP phases in the mixing matrix. Indeed, CP phases introduced in the nine Euler-angle-like parameterizations Fritzsch:1997st appear to take unrelated and scattered values at first glance.
Recently, some studies have begun focusing on rephasing invariants that explicitly incorporate the determinant of the mixing matrix $V$ Yang:2025hex . A specific relation $\delta_{\rm PDG}+\delta_{\rm KM}=\pi-\alpha+\gamma$ Yang:2025ftl between CP phases and the angles and its generalized sum rules Yang:2025vrs reveal an underlying regularity between chaotic phases and angles. Historically, in particle physics, such regularities have often indicated the existence of more fundamental physical quantities.
In this study, we present six third-order rephasing invariants with determinant Yang:2025law ; Luo:2025wio reproduce all nine CP angles of the unitarity triangles and all CP phases in the nine different parameterizations. The sum rules relating the CP angles and phases are also decomposed into expressions of these fewer invariants. Furthermore, through ninth-order invariants, we demonstrate the existence of a duality between the nine CP phases and the nine angles.
@@ II Nine angles of unitarity triangles and nine CP phases in parameterizations of mixing matrix
Here, we first review the general sum rules between nine angles of the unitarity triangles and CP phases in the nine Euler-angle-like parameterizations of the mixing matrix Fritzsch:1997st . From the latest UTfit values for the CKM matrix UTfit:2022hsi ,
$\sin\theta_{12}$ $=0.22519\pm 0.00083\,,~~~\sin\theta_{23}=0.04200\pm 0.00047\,$, (1)
$\sin\theta_{13}$ $=0.003714\pm 0.000092\,,~~~\delta=1.137\pm 0.022=65.15^{\circ}\pm 1.3^{\circ}\,.$ (2)
The angles of the unitarity triangles from the best-fit values are
@@
where $\Pi_{\alpha i}:=V_{\beta j}V_{\beta k}^{*}V_{\gamma k}V_{\gamma j}^{*}$ and
$\begin{pmatrix}V_{tb}V_{ts}^{*}V_{cs}V_{cb}^{*}&V_{td}V_{tb}^{*}V_{cb}V_{cd}^{*}&V_{ts}V_{td}^{*}V_{cd}V_{cs}^{*}\\ V_{ub}V_{us}^{*}V_{ts}V_{tb}^{*}&V_{ud}V_{ub}^{*}V_{tb}V_{td}^{*}&V_{us}V_{ud}^{*}V_{td}V_{ts}^{*}\\ V_{cb}V_{cs}^{*}V_{us}V_{ub}^{*}&V_{cd}V_{cb}^{*}V_{ub}V_{ud}^{*}&V_{cs}V_{cd}^{*}V_{ud}V_{us}^{*}\end{pmatrix}$. (5)
@@
On the other hand, Fritzsch and Xing have pointed out that there exist nine inequivalent Euler-angle parameterizations of the mixing matrix Fritzsch:1997st . Since these matrix elements have trivial phases in certain row $\alpha$ and column $i$ with $V_{\alpha j},V_{\beta i}\in\mathbb{R}$, we denote their phases by $\delta^{(\alpha i)}$ to distinguish between them.
In a rephasing invariant manner, these phases $\delta^{(\alpha i)}$ are evaluated as Yang:2025vrs
$\delta^{(11)}$ $=\arg\left[{V_{12}V_{13}V_{21}V_{31}\over V_{11}\det V}\right],~~\delta^{(12)}=\arg\left[{V_{11}V_{13}V_{22}V_{32}\over V_{12}\det V}\right],~~\delta^{(13)}=\arg\left[{V_{11}V_{12}V_{23}V_{33}\over V_{13}\det V}\right],$ (7)
$\delta^{(21)}$ $=\arg\left[{V_{22}V_{23}V_{11}V_{31}\over V_{21}\det V}\right],~~\delta^{(22)}=\arg\left[{V_{21}V_{23}V_{12}V_{32}\over V_{22}\det V}\right],~~\delta^{(23)}=\arg\left[{V_{21}V_{22}V_{13}V_{33}\over V_{23}\det V}\right],$ (8)
@@
From such expressions, it is not difficult to imagine that nontrivial relations exist between $\Phi_{\alpha i}$ and $\delta^{(\alpha i)}$. In fact, the following exact sum rule between the phases and the angles has been discovered Yang:2025ftl
@@
where the phase matrix $\Delta$ and the transfer matrix $T$ are
@@
Many of these nontrivial relations suggest the existence of more fundamental underlying quantities. In the next section, we show that all nine CP angles and all nine CP phases are derived by six third-order rephasing invariants.
@@ III Third-order rephasing invariants with determinant
In order to express these CP-violating observables by fewer invariants, we employ the six rephasing invariants as follows Yang:2025law ; Luo:2025wio .
$\chi_{1}$ $=\arg\left[\frac{V_{11}V_{22}V_{33}}{\det V_{\rm CKM}}\right]=-0.0019^{\circ}\,,~\chi_{2}=\arg\left[\frac{V_{12}V_{23}V_{31}}{\det V_{\rm CKM}}\right]=-22.45^{\circ},~\chi_{3}=\arg\left[\frac{V_{13}V_{21}V_{32}}{\det V_{\rm CKM}}\right]=-64.06^{\circ}\,$,
$\psi_{1}$ $=\arg\left[-\frac{V_{11}V_{23}V_{32}}{\det V_{\rm CKM}}\right]=1.05^{\circ}\,,~\psi_{2}=\arg\left[-\frac{V_{12}V_{21}V_{33}}{\det V_{\rm CKM}}\right]=0.035^{\circ}\,,~\psi_{3}=\arg\left[-\frac{V_{13}V_{22}V_{31}}{\det V_{\rm CKM}}\right]=92.40^{\circ}\,.$
@@
$\psi_{1}+\psi_{2}+\psi_{3}=\arg\left[-\frac{\prod_{\alpha,i}V_{\alpha i}}{(\det V_{\rm CKM})^{3}}\right]=\chi_{1}+\chi_{2}+\chi_{3}+\pi=93.49^{\circ}\,.$ (14)
@@
Since the second column corresponds to the unitarity triangle angles, $\psi_{2}-\chi_{2}=\beta\,,\psi_{3}-\chi_{1}=\alpha\,,\psi_{1}-\chi_{3}=\gamma$ hold. Moreover, since the three invariants have arguments smaller than about $1^{\circ}$, the following notation can also be meaningful:
@@
$\psi_{1}=d\gamma\,,~~\psi_{2}$ $=d\beta\,,~~\psi_{3}=\alpha-d\alpha\,.$ (17)
Then, the nine CP angles are
@@
and the sum of each row and column clearly reduces to Eq. (6), $\alpha+\beta+\gamma=\pi$.
The three small quantities in the standard PDG parameterization are
$\chi_{1}\simeq-s_{12}s_{13}s_{23}\sin\delta\sim\lambda^{6}\,,~~\psi_{2}\simeq\frac{s_{13}s_{23}}{s_{12}}\sin\delta\sim\lambda^{4}\,,~~\psi_{1}\simeq\frac{s_{12}s_{13}}{s_{23}}\sin\delta\sim\lambda^{2}\,$, (19)
@@
It is evident that the phase matrix $\Delta_{ij}=\delta^{(ij)}$ can be expressed in terms of the six invariants and the nine CP angles. To obtain a representation of $\Delta$ in which $\pi$ does not explicitly appear, we can only use $\Phi$ and $\Psi$. Furthermore, since $\det V$ appears in the denominator of $\Delta_{ij}$, the only possibility is to add $\Phi_{ij}$ and $\psi_{i}$. Among these sums, there are two cases in which the denominator $V_{ij}$ does not overlap, allowing them to be written as two matrix equations.
@@
$\Delta$ $=\begin{pmatrix}-\chi_{1}+\psi_{2}+\psi_{3}&-\chi_{2}+\psi_{3}+\psi_{1}&-\chi_{3}+\psi_{1}+\psi_{2}\\ -\chi_{3}+\psi_{3}+\psi_{1}&-\chi_{1}+\psi_{1}+\psi_{2}&-\chi_{2}+\psi_{2}+\psi_{3}\\ -\chi_{2}+\psi_{1}+\psi_{2}&-\chi_{3}+\psi_{2}+\psi_{3}&-\chi_{1}+\psi_{3}+\psi_{1}\\ \end{pmatrix}$ (23)
@@
From these definitions, sum rules between the CP phases and angles are described by more fundamental invariants. First, let us examine how the specific sum rule (11) is expressed in terms of invariants $\psi_{i}$ and $\chi_{i}$. From Eqs. (16) and (23),
@@
$\begin{pmatrix}\psi_{1}&\psi_{2}&\psi_{3}\\ \psi_{2}&\psi_{3}&\psi_{1}\\ \psi_{3}&\psi_{1}&\psi_{2}\\ \end{pmatrix}$ $=\begin{pmatrix}\delta^{(12)}&\delta^{(13)}&\delta^{(11)}\\ \delta^{(22)}&\delta^{(23)}&\delta^{(21)}\\ \delta^{(32)}&\delta^{(33)}&\delta^{(31)}\end{pmatrix}-\begin{pmatrix}\Phi_{31}&\Phi_{32}&\Phi_{33}\\ \Phi_{11}&\Phi_{12}&\Phi_{13}\\ \Phi_{21}&\Phi_{22}&\Phi_{23}\\ \end{pmatrix}$ (32)
@@
the essence of the sum rules is that they constitute an algebraically trivial identity. Eq. (29) corresponds to the 1-2 element of Eqs. (32) and (33).
@@
However, it corresponds to the same conditions as Eqs. (32) and (33), and does not introduce any additional constraints. This is because Eqs. (31) and (35) are equivalent under $\Psi=T\Psi T$.
Such relations can also be expressed by $\rm X$} instead of $\Psi$. Let us note the following fact.
@@
Defining $\Pi\equiv(1+T+T^{2})\pi$ with $\Pi_{ij}=\pi$, it follows from Eq. (6) that
$(1+T+T^{2})\Phi=\Pi\,.$ (37)
By substituting this together with $\Phi=\Psi-X$ into Eq. (30), it is evident that
$\Delta T=\Psi+T\Phi={\rm X}+\Pi-T^{T}\Phi\,,~\Delta T^{T}=\Psi+T^{T}\Phi={\rm X}+\Pi-T\Phi\,.$ (38)
Eliminating $X+\Pi$ reduces the expression (12). On the other hand, solving for $X+\Pi$ yields
${\rm X}+\Pi=\Delta T+T^{T}\Phi=\Delta T^{T}+T\Phi\,.$ (39)
@@
$\begin{pmatrix}\chi_{1}&\chi_{2}&\chi_{3}\\ \chi_{3}&\chi_{1}&\chi_{2}\\ \chi_{2}&\chi_{3}&\chi_{1}\\ \end{pmatrix}+\begin{pmatrix}\pi&\pi&\pi\\ \pi&\pi&\pi\\ \pi&\pi&\pi\\ \end{pmatrix}$ $=\begin{pmatrix}\delta^{(12)}&\delta^{(13)}&\delta^{(11)}\\ \delta^{(22)}&\delta^{(23)}&\delta^{(21)}\\ \delta^{(32)}&\delta^{(33)}&\delta^{(31)}\end{pmatrix}+\begin{pmatrix}\Phi_{21}&\Phi_{22}&\Phi_{23}\\ \Phi_{31}&\Phi_{32}&\Phi_{33}\\ \Phi_{11}&\Phi_{12}&\Phi_{13}\\ \end{pmatrix}$ (40)
@@
$\arg\left[-\frac{V_{12}V_{23}V_{31}}{\det V_{\rm CKM}}\right]=\arg\left[{V_{11}V_{12}V_{23}V_{33}\over V_{13}\det V_{\rm CKM}}\right]+\arg\left[-{V_{13}V_{31}\over V_{11}V_{33}}\right]=\arg\left[{V_{12}V_{13}V_{21}V_{31}\over V_{11}\det V_{\rm CKM}}\right]+\arg\left[-{V_{23}V_{11}\over V_{21}V_{13}}\right].$ (43)
The reason is that the fundamental equations consist of Eqs. (16) and (21), and the rewriting from $\Psi$ into X relies on Eqs. (16), $\Psi={\rm X}+\Phi$. While the representation is doubled with respect to $\Psi$ and X, the sum rules for $\Phi$ and $\Delta$ stay at the same. On the other hand, since there are four ways to construct a third-order invariant from products of a fifth-order invariant and fourth-order invariants, such representations exist in 36 forms for the nine fifth-order invariants, and the counting is consistent.
@@
$T\Delta=\Psi+\Phi T=\Pi+{\rm X}-\Phi T^{T}\,,~~T^{T}\Delta=\Psi+\Phi T^{T}=\Pi+{\rm X}-\Phi T\,.$ (45)
As anticipated, these conditions are also not independent. Since these are decompositions of the sum rules into the six fundamental invariants, it is free of redundancy and easier to handle.

### v2

@@ Abstract
In this paper, we demonstrate that products of third-order rephasing invariants $V_{\alpha i}V_{\beta j}V_{\gamma k}/\det V$ of flavor mixing matrix $V$ reproduce all the nine angles of unitarity triangles and all the CP phases in the nine parameterizations of $V$. The sum rules relating the CP phases and angles are also decomposed into terms of these rephasing invariants. Furthermore, through ninth-order invariants, these fourth- and fifth-order invariants become equivalent, which can be regarded as a certain duality. For the phase matrix $\Delta$ and the angle matrix $\Phi$, $\Delta\pm\Phi$ are expressed in terms of even-permutations X and odd-permutations $\Psi$ of third-order invariants. As a result, these are represented by the two concise matrix equations $\Phi=\Psi-{\rm X}$ and $\Delta=\Pi^{\prime}-\Psi-{\rm X}$.
@@ I Introduction
The flavor mixing matrix provides an essential formalism of flavor physics. Since its CP phase is intimately related to the origin of the matter-antimatter asymmetry, it is one of the key observables for particle physics in the near future. In order to establish a theoretical basis, rephasing invariants of the mixing matrix Wu:1985ea ; Bernabeu:1986fc ; Gronau:1986xb ; Branco:1987mj ; Bjorken:1987tr ; Nieves:1987pp ; Botella:1994cs ; Branco:1999fs ; Jenkins:2007ip have been studied to define CP-violating quantities and to elucidate their physical significance. The fourth-order rephasing invariants, represented by the Jarlskog invariant Jarlskog:1985ht , have been widely employed in the analysis of unitarity triangles Wu:1994di ; Hocker:2006xb ; Xing:2009eg ; Harrison:2009bz ; Antusch:2009hq ; Frampton:2010ii ; Dueck:2010fa ; Frampton:2010uq ; Li:2010ae ; Qin:2010hn ; Zhou:2011xm ; Qin:2011ub ; Zhang:2012bk ; Zhang:2012ys ; Li:2012zxa ; He:2013rba ; He:2016dco ; Xing:2019tsn ; Kaur:2023ypg ; Harrison:2025rkp . However, they are insufficient for providing a unified understanding of the CP phases in the mixing matrix. Indeed, CP phases introduced in the nine Euler-angle-like parameterizations Fritzsch:1997st appear to take unrelated and scattered values at first glance.
Recently, some studies have begun focusing on rephasing invariants that explicitly incorporate the determinant of the mixing matrix $V$ Yang:2025hex . Unlike the Jarlskog-type quartic invariants, these new invariants involving $\det V$ provide a concise framework that can express the CP-violating phases. A specific relation $\delta_{\rm PDG}+\delta_{\rm KM}=\pi-\alpha+\gamma$ presented a nontrivial connection between the CP phases $\delta_{\rm PDG},\delta_{\rm KM}$ and the angles $\alpha,\gamma$ of the unitarity triangle Yang:2025ftl . Furthermore, the set of nine formulae for the CP phases $\delta^{(\alpha i)}=\arg[{V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}/V_{\alpha i}^{3}\det V}]$ and its generalized sum rules Yang:2025vrs reveal an underlying regularity between chaotic CP phases and unitarity triangles. Historically, in particle physics, such regularities have often indicated the existence of more fundamental physical quantities.
In this study, we present six third-order rephasing invariants with determinant Kuo:2005pf ; Yang:2025law ; Luo:2025wio reproduce all nine CP angles of the unitarity triangles and all CP phases in the nine different parameterizations. The sum rules relating the CP angles and phases are also decomposed into expressions of these fewer invariants. Furthermore, through ninth-order invariants, we demonstrate the existence of a duality between the nine CP phases and the nine angles.
This paper is organized as follows. The next section gives a rephasing invariant formalism of unitarity triangles and CP phases of the mixing matrix. In Sec. 3, we demonstrate decomposition of sum rules by third-order rephasing invariants and a duality between the CP phases and the unitarity triangles. The final section is devoted to a summary.
@@ II Nine angles of unitarity triangles and nine CP phases in parameterizations of mixing matrix
Here, we first review the general sum rules between nine angles of the unitarity triangles and CP phases in the nine Euler-angle-like parameterizations of the mixing matrix. The observed values of the CKM matrix from the latest UTfit are UTfit:2022hsi ,
$\sin\theta_{12}$ $=0.22519\pm 0.00083\,,~~~\sin\theta_{23}=0.04200\pm 0.00047\,$,
$\sin\theta_{13}$ $=0.003714\pm 0.000092\,,~~~\delta=1.137\pm 0.022=65.15^{\circ}\pm 1.3^{\circ}\,$, (1)
and the angles of the unitarity triangles for the best-fit values are
@@
where $\Pi_{\alpha i}\equiv V_{\beta j}V_{\beta k}^{*}V_{\gamma k}V_{\gamma j}^{*}$ and
$\Pi=\begin{pmatrix}V_{tb}V_{ts}^{*}V_{cs}V_{cb}^{*}&V_{td}V_{tb}^{*}V_{cb}V_{cd}^{*}&V_{ts}V_{td}^{*}V_{cd}V_{cs}^{*}\\ V_{ub}V_{us}^{*}V_{ts}V_{tb}^{*}&V_{ud}V_{ub}^{*}V_{tb}V_{td}^{*}&V_{us}V_{ud}^{*}V_{td}V_{ts}^{*}\\ V_{cb}V_{cs}^{*}V_{us}V_{ub}^{*}&V_{cd}V_{cb}^{*}V_{ub}V_{ud}^{*}&V_{cs}V_{cd}^{*}V_{ud}V_{us}^{*}\end{pmatrix}$. (4)
@@
On the other hand, Fritzsch and Xing have pointed out that there exist nine inequivalent Euler-angle parameterizations of the mixing matrix Fritzsch:1997st . Since these matrix elements have trivial phases in certain row $\alpha$ and column $i$ with $V_{\alpha j},V_{\beta i}\in\mathbb{R}$, we denote their phases by $\delta^{(\alpha i)}$ to distinguish between them. The structure of CP phases in such a parameterization $V^{(\alpha i)}$ is determined by the following conditions
$\arg\det V^{(\alpha i)}=-\delta^{(\alpha i)}\,,~\arg V_{\alpha 1}^{(\alpha i)}=\arg V_{\alpha 2}^{(\alpha i)}=\arg V_{\alpha 3}^{(\alpha i)}=\arg V_{1i}^{(\alpha i)}=\arg V_{2i}^{(\alpha i)}=\arg V_{3i}^{(\alpha i)}=0~{\rm or}~\pi\,.$ (6)
Since there is a duplication for the element $V_{\alpha i}^{(\alpha i)}$, there are effectively six independent conditions, and the phase $\pi$ appears either zero or two times.
To determine these phases $\delta^{(\alpha i)}$, we consider rephasing transformations from $V^{(\alpha i)}$ to the mixing matrix $V$ in an arbitrary basis
$V=\Psi_{L}V^{(\alpha i)}\Psi_{R}^{\dagger}\,,~~V_{\beta j}=e^{i\gamma_{L\beta}}V^{(\alpha i)}_{\beta j}e^{-i\gamma_{Rj}}\,$, (7)
where $(\Psi_{L})_{\alpha\beta}=e^{i\gamma_{L\alpha}}\delta_{\alpha\beta}$ and $(\Psi_{R})_{ij}=e^{i\gamma_{Ri}}\delta_{ij}$ are diagonal phase matrices. Comparing the determinants of the two representations, one finds $\arg\det V=-\delta^{(\alpha i)}+\sum_{\beta,j}(\gamma_{L\beta}-\gamma_{Rj})$ and the CP phases $\delta^{(\alpha i)}$ Yang:2025vrs
$\delta^{(\alpha i)}$ $=(\gamma_{L1}+\gamma_{L2}+\gamma_{L3}-\gamma_{R1}-\gamma_{R2}-\gamma_{R3})-\arg\det V$
$=\arg[V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}/V_{\alpha i}^{3}]-\arg\det V=\arg\left[{V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}\over V_{\alpha i}^{3}\det V}\right].$ (8)
More specifically, the two factors $V_{\alpha i}$ in the numerator cancel those in the denominator,
$\delta^{(11)}$ $=\arg\left[{V_{12}V_{13}V_{21}V_{31}\over V_{11}\det V}\right],~~\delta^{(12)}=\arg\left[{V_{11}V_{13}V_{22}V_{32}\over V_{12}\det V}\right],~~\delta^{(13)}=\arg\left[{V_{11}V_{12}V_{23}V_{33}\over V_{13}\det V}\right],$
$\delta^{(21)}$ $=\arg\left[{V_{22}V_{23}V_{11}V_{31}\over V_{21}\det V}\right],~~\delta^{(22)}=\arg\left[{V_{21}V_{23}V_{12}V_{32}\over V_{22}\det V}\right],~~\delta^{(23)}=\arg\left[{V_{21}V_{22}V_{13}V_{33}\over V_{23}\det V}\right],$
@@
From such expressions, some nontrivial relations are expected between $\Phi_{\alpha i}$ and $\delta^{(\alpha i)}$. In fact, the following exact sum rule between the phases and the angles has been discovered Yang:2025ftl
@@
where the phase matrix $\Delta$ and the transfer matrix $T$ are defined as
@@
Many of these nontrivial relations suggest the existence of more fundamental underlying quantities. In the next section, we show that all nine CP angles and all nine CP phases are derived by third-order rephasing invariants.
@@ III Third-order rephasing invariants with determinant
In order to express these CP-violating observables by fewer invariants, we employ the six rephasing invariants as follows Kuo:2005pf ; Yang:2025law ; Luo:2025wio .
$\chi_{1}$ $=\arg\left[\frac{V_{11}V_{22}V_{33}}{\det V}\right]=-0.0019^{\circ}\,,~\chi_{2}=\arg\left[\frac{V_{12}V_{23}V_{31}}{\det V}\right]=-22.45^{\circ},~\chi_{3}=\arg\left[\frac{V_{13}V_{21}V_{32}}{\det V}\right]=-64.06^{\circ}\,$,
$\psi_{1}$ $=\arg\left[-\frac{V_{11}V_{23}V_{32}}{\det V}\right]=1.05^{\circ}\,,~\psi_{2}=\arg\left[-\frac{V_{12}V_{21}V_{33}}{\det V}\right]=0.035^{\circ}\,,~\psi_{3}=\arg\left[-\frac{V_{13}V_{22}V_{31}}{\det V}\right]=92.40^{\circ}\,.$
The arguments of these third-order invariants correspond to the nontrivial phases of the matrix elements $V^{0}_{\alpha i}$ in the PDG parametrization;
$\chi_{1}=\arg V_{22}^{0}\,,~~\chi_{2}=\arg V_{31}^{0}\,$,~ $~\chi_{3}=\arg V_{13}^{0}+\arg V_{21}^{0}+\arg V_{32}^{0}\,$,~~
$\psi_{1}=\arg V_{32}^{0}+\pi\,,~~\psi_{2}=\arg V_{21}^{0}+\pi\,$,~ $~\psi_{3}=\arg V_{13}^{0}+\arg V_{22}^{0}+\arg V_{31}^{0}+\pi\,.$ (14)
@@
$\psi_{1}+\psi_{2}+\psi_{3}=\arg\left[-\frac{\prod_{\alpha,i}V_{\alpha i}}{(\det V)^{3}}\right]=\chi_{1}+\chi_{2}+\chi_{3}+\pi=93.49^{\circ}\,.$ (15)
@@
The subsequent letter Yang:2025dhm shows that these arguments of the invariants $\Psi$, $\rm X$}, and $\Phi$ correspond to the angles of an alternative set of unitarity triangles obtained from the inversion formula $U^{\dagger}=U^{-1}$. Therefore, the sum rules between CP phases and angles presented below are expected to be understood geometrically by compositions of the new unitarity triangles.
Since the second column corresponds to the unitarity triangle angles, $\psi_{2}-\chi_{2}=\beta\,,\psi_{3}-\chi_{1}=\alpha\,,\psi_{1}-\chi_{3}=\gamma$ hold. Moreover, since the three angles $\chi_{1}\,,\psi_{1}\,,\psi_{2}$ are smaller than about $1^{\circ}$, it is also useful to regard these angles as deviations associated with $\alpha,\beta,\gamma$,
@@
$\psi_{1}=d\gamma\,,~~\psi_{2}$ $=d\beta\,,~~\psi_{3}=\alpha-d\alpha\,$, (18)
where $d$ symbolically denotes small differences. Then, the nine CP angles are
@@
and the sum of each row and column clearly reduces to Eq. (5), $\alpha+\beta+\gamma=\pi$. The three small quantities in the standard PDG parameterization are
$\chi_{1}\simeq-s_{12}s_{13}s_{23}\sin\delta\sim\lambda^{6}\,,~~\psi_{2}\simeq\frac{s_{13}s_{23}}{s_{12}}\sin\delta\sim\lambda^{4}\,,~~\psi_{1}\simeq\frac{s_{12}s_{13}}{s_{23}}\sin\delta\sim\lambda^{2}\,.$ (20)
@@
It is evident that the phase matrix $\Delta_{\alpha i}=\delta^{(\alpha i)}$ can be expressed in terms of the six invariants and the nine CP angles. To obtain a representation of $\Delta$ in which $\pi$ does not explicitly appear, we can only use $\Phi$ and $\Psi$. Furthermore, since $\det V$ appears in the denominator of $\Delta_{\alpha i}$, the only possibility is to add $\Phi_{\alpha i}$ and $\psi_{i}$. Among the three possibilities of $\psi_{i}$, there are two cases in which the denominator of $\Delta_{\alpha i}$ does not overlap, allowing them to be written as two matrix equations.
@@
$\Delta$ $=\begin{pmatrix}-\chi_{1}+\psi_{2}+\psi_{3}&-\chi_{2}+\psi_{3}+\psi_{1}&-\chi_{3}+\psi_{1}+\psi_{2}\\ -\chi_{3}+\psi_{3}+\psi_{1}&-\chi_{1}+\psi_{1}+\psi_{2}&-\chi_{2}+\psi_{2}+\psi_{3}\\ -\chi_{2}+\psi_{1}+\psi_{2}&-\chi_{3}+\psi_{2}+\psi_{3}&-\chi_{1}+\psi_{3}+\psi_{1}\\ \end{pmatrix}$
@@
Thus, although these fifth-order invariants are apparently irreducible, they are expressed as a product of three third-order invariants.
@@
From these definitions, sum rules between the CP phases and angles are described by more fundamental invariants. First, let us examine how the specific sum rule (11) is expressed in terms of invariants $\psi_{i}$ and $\chi_{i}$. From Eqs. (17) and (III),
@@
$\begin{pmatrix}\psi_{1}&\psi_{2}&\psi_{3}\\ \psi_{2}&\psi_{3}&\psi_{1}\\ \psi_{3}&\psi_{1}&\psi_{2}\\ \end{pmatrix}$ $=\begin{pmatrix}\delta^{(12)}&\delta^{(13)}&\delta^{(11)}\\ \delta^{(22)}&\delta^{(23)}&\delta^{(21)}\\ \delta^{(32)}&\delta^{(33)}&\delta^{(31)}\end{pmatrix}-\begin{pmatrix}\Phi_{31}&\Phi_{32}&\Phi_{33}\\ \Phi_{11}&\Phi_{12}&\Phi_{13}\\ \Phi_{21}&\Phi_{22}&\Phi_{23}\\ \end{pmatrix}$
@@
the essence of the sum rules is that they constitute an algebraically trivial identity for $\chi_{i}$ and $\psi_{i}$. Eq. (29) corresponds to the 1-2 element of Eq. (32).
@@
However, it corresponds to the same conditions as Eq. (32), and does not introduce any additional constraints. This is because Eqs. (31) and (34) are equivalent under $\Psi=T\Psi T$.
Such relations can also be expressed by $\rm X$} instead of $\Psi$. Let us note the following fact
@@
From this fact, Eq. (5) is rewritten as
$(1+T+T^{2})\Phi=(1+T+T^{2})\pi\,.$ (36)
By substituting this together with $\Phi=\Psi-{\rm X}$ into Eq. (30), it is evident that
$\Delta T=\Psi+T\Phi={\rm X}+(1+T+T^{2})\pi-T^{T}\Phi\,,~\Delta T^{T}=\Psi+T^{T}\Phi={\rm X}+(1+T+T^{2})\pi-T\Phi\,.$ (37)
Eliminating $\rm X$} reduces the expression (12). On the other hand, solving for $\rm X$} yields
${\rm X}+(1+T+T^{2})\pi=\Delta T+T^{T}\Phi=\Delta T^{T}+T\Phi\,.$ (38)
@@
$\begin{pmatrix}\chi_{1}&\chi_{2}&\chi_{3}\\ \chi_{3}&\chi_{1}&\chi_{2}\\ \chi_{2}&\chi_{3}&\chi_{1}\\ \end{pmatrix}+\begin{pmatrix}\pi&\pi&\pi\\ \pi&\pi&\pi\\ \pi&\pi&\pi\\ \end{pmatrix}$ $=\begin{pmatrix}\delta^{(12)}&\delta^{(13)}&\delta^{(11)}\\ \delta^{(22)}&\delta^{(23)}&\delta^{(21)}\\ \delta^{(32)}&\delta^{(33)}&\delta^{(31)}\end{pmatrix}+\begin{pmatrix}\Phi_{21}&\Phi_{22}&\Phi_{23}\\ \Phi_{31}&\Phi_{32}&\Phi_{33}\\ \Phi_{11}&\Phi_{12}&\Phi_{13}\\ \end{pmatrix}$
@@
$\arg\left[-\frac{V_{12}V_{23}V_{31}}{\det V}\right]=\arg\left[{V_{11}V_{12}V_{23}V_{33}\over V_{13}\det V}\right]+\arg\left[-{V_{13}V_{31}\over V_{11}V_{33}}\right]=\arg\left[{V_{12}V_{13}V_{21}V_{31}\over V_{11}\det V}\right]+\arg\left[-{V_{23}V_{11}\over V_{21}V_{13}}\right].$ (41)
The reason is that the fundamental equations consist of Eqs. (17) and (22), and the rewriting from $\Psi$ into X relies on Eqs. (17), $\Psi={\rm X}+\Phi$. While the representation is doubled with respect to $\Psi$ and X, the sum rules for $\Phi$ and $\Delta$ stay the same. On the other hand, since there are four ways to construct a third-order invariant from products of a fifth-order invariant and fourth-order invariants, such representations exist in 36 forms for the nine fifth-order invariants, and the counting is consistent.
@@
$T\Delta=\Psi+\Phi T=(1+T+T^{2})\pi+{\rm X}-\Phi T^{T}\,,~~T^{T}\Delta=\Psi+\Phi T^{T}=(1+T+T^{2})\pi+{\rm X}-\Phi T\,.$ (43)
As anticipated, these conditions are also not independent. Since these are decompositions of the sum rules into the six fundamental invariants, it will be free of redundancy and easier to handle.